The area of the United States is about 3,600,000 square miles. What percent of the Earth's surface is the area of the USA? The radius of the Earth is about 3950 miles on average. Rond answer to the nearest tenth.
Answer by checkley77(12844)

TSA=4PIR^2
TSA=4*3.14*3,950^2
TSA=12.56*15,602,500
TSA=195,967,400 TOTAL AREA OF THE EARTH.
3,600,000/195,967,400=.01837 OR 1.84 %.